the resistance of a 31.4m long wire is 20ohms. find the resistivity of the wire if it has 1mm dia
 

Grade:10

1 Answers

Beenu Mathew
7 years ago
HI,
 
ITS A DIRECT QUESTION
LOOK
R=20 OHMS
X= RESISTIVITY (TO BE FOUND OUT)
l= 31.4 m or ( pie * 10)
A = pie d2/4  (where A-area....n d...is diameter)
R=Xl/A
RA/l =X
X= 20 * pie (10-3)2/4 *pie * 10 
on solving this ….u will get
X= 0.5 * 10-6 ohm m (OR) 5* 10-7 ohm m
 
JEHOYESJIRA